A rectangular garden has dimensions of 18 feet by 13 feet. A gravel path of equal width is to be built around the garden. How wide can the path be if there is enough gravel for 516 square feet?
Answer by checkley77(12844)

18*13=234 FT^2 FOR THE GARDEN.
234+516=750 FOR THE TOTAL AREA WITH THE PATH.
(18+2X)(13+2X)=750
234+26X+36X+4X^2=750
4X^2+62X+234-750=0
4X^2+62X-516=0
2(2X^2+31X-258)=0
2(2X-12)(X+21.5)=0
2X-12=0
2X=12
X=6 FT. IS THE ANSWER FOR THE WIDTH OF THE PATH.
PROOF:
(18+2*6)(13+2*6)=750
(18+12)(13+12)=750
30*25=750
750=750